After a nice summer visit, you are taking your friend to the Train Station to return home. The Train Station is 90 miles from your home. You drive to get some fast food which is 5 miles from your home on the way to the Train Station. After lunch you leave the fast food restaurant at 1:00 PM and travel at a constant rate of 40 mph to the Train Station. The Train departs at 3:00 PM.


1
Expert's answer
2021-06-17T16:03:57-0400

After getting fast food, 5 miles is already covered

Remaining distance = (90 - 5) miles =85 miles

Time taken to the train station = Distance/speed = 85/40 =2.125 hours

This implies that we will reach the train station at 1.00 pm +2 hrs 7.5 minutes = 3. 08 pm

Since the train departs at 3.00 pm, this means that we will not reach the train station on time, and will thus miss the train.